OpenStudy (anonymous):

Write the standard form of the equation of each line passing through P with slope m. P=(-5,2); m=1/3

OpenStudy (marissalovescats):

I think there is an equation for this but I don't believe I recall it. Let me take a gander at google search :P

OpenStudy (anonymous):

is this an exercise for someone?

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (marissalovescats):

y−y1=m(x−x1) thats the one i think you should use

OpenStudy (anonymous):

ok i am totally new to this stuff so any help in explaining this stuff small steps at a time would be helpful. thanks

OpenStudy (marissalovescats):

Because you have a point and the slope. So like y-2=1/3(x+5)

OpenStudy (jhannybean):

Just use the slope intercept form to find the value of the y-intercept, then write your equation out in slope intercept form.

OpenStudy (marissalovescats):

How I put it into the equation: